Each question on a multiple choice test has five equally likely answers. If Nikka guesses randomly on four questions, what is the probability that she will answer at least one correctly?

Answer by stanbon(75887)

Each question on a multiple choice test has five equally likely answers. If Nikka guesses randomly on four questions, what is the probability that she will answer at least one correctly?
----
Binomial Problem with n = 4 and p(correct) = 1/5 and P(wrong) = 4/5
----------------------------
P(at least one correct)
= 1 - P(none correct)
----
= 1 - P(all wrong)
----
= 1 - (4/5)^4
----
= 1 - 0.4096
-----
= 0.5904
==============
Cheers,
Stan H.